# Landauer’s principle and black hole area quantization

### v2

@@ Abstract
$\frac{S_{\rm BH}}{k_{B}}=\frac{c^{3}}{4\hbar G}A=\frac{\pi G}{c\hbar}M^{2},$ (1)
where $k_{B}$ is the Boltzmann constant and $\hbar$ is the (reduced) Planck’s constant. It clearly reveals that the horizon area of the black hole also increases irreversibly when matter is absorbed, with $A$ serving as a measure of information bek2 . The corresponding temperature, namely, the Hawking temperature, can be easily recognized to be
@@
Finally, let us discuss the effect of the inclusion of angular momentum as in the Kerr black hole whose mass reads as
@@
where $0\leq a\leq 1$ is the rotation parameter and we have taken $G=l_{P}^{2}=1$ so that our expressions match with those of 126 . Notice that upon putting $a=0$ we get $T_{\rm H}=1/(8\pi M)$ and $\Omega_{\rm H}=0$, corresponding to the Schwarzschild black hole. Eq. (19) lets us write $\Omega_{\rm H}=\frac{2\pi T_{\rm H}a}{\sqrt{1-a^{2}}}$, implying that

### v3

@@ Abstract
$\frac{S_{\rm BH}}{k_{B}}=\frac{c^{3}}{4\hbar G}A=\frac{4\pi G}{c\hbar}M^{2},$ (1)
where $k_{B}$ is the Boltzmann constant and $\hbar$ is the (reduced) Planck’s constant. It clearly reveals that the horizon area of the black hole also increases irreversibly when matter is absorbed with $A$ serving as a measure of information bek2 . The corresponding temperature, namely, the Hawking temperature, can be easily recognized to be
@@
Finally, let us discuss the effect of the inclusion of angular momentum as in the Kerr black hole whose mass reads as (we take $G=l_{P}^{2}=1$ so that our equations are consistent with those of 126 )
@@
where $0\leq a\leq 1$ is the rotation parameter. Notice that upon putting $a=0$ we get $T_{\rm H}=1/(8\pi M)$ and $\Omega_{\rm H}=0$, corresponding to the Schwarzschild black hole. Eq. (19) lets us write $\Omega_{\rm H}=\frac{2\pi T_{\rm H}a}{\sqrt{1-a^{2}}}$, implying that